OpenStudy (danysanchy97):

Lavina wants to buy a rocking chair for $160. She will pay 10% down and pay the rest in six monthly installments. What will be the amount of each monthly payment? A. $27 B. $26 C. $24 D. $16

OpenStudy (danysanchy97):

would this be B?

OpenStudy (danysanchy97):

@KamiBug @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

each month 10% of that cost is paid off

OpenStudy (kamibug):

I think that 10% is going to be paid off first, because that's the down payment. Then you gotta divide what's left between 6 months. So... 0.10 * 160 = 16 160 - 16 = 144 144 / 6 = ?

OpenStudy (danysanchy97):

thank you